OpenStudy (anonymous):

The sum of the measures of two complementary angles exceeds the difference of their measures by 72°. Find the measure of the smaller angle. A. 18° B. 36° C. 54° D. 58°

OpenStudy (anonymous):

x + y = 90 (complimentary) 90 - 72 = x - y 18 = x - y Solve: 18 = x - y x + y = 90

OpenStudy (anonymous):

I think its A